# Dislocation-mediated short-range order evolution during thermomechanical processing

### v1

@@ 2.4. SRO quantification
To quantify chemical SRO in the states obtained from both TMP and annealing simulations, we employed the chemical motif framework introduced in refs. [14] and [15]. In this approach, the local chemical environment of an atom was described by a motif $\mathcal{M}_{i}$, consisting of the atom and its first-nearest neighbors. We utilized the Euclidean graph neural network architecture[62] from ref. [15] to identify all motifs $\mathcal{M}_{i}$ in a system and compute their population density distributions $P(\mathcal{M}_{i})$, hereafter denoted simply as $P$ for brevity. The difference in SRO between two states, characterized by motif population density distributions $P_{1}$ and $P_{2}$, was quantified using the Jensen-Shannon divergence:
@@
where $M=\frac{1}{2}(P_{1}+P_{2})$ and $D_{\text{KL}}$ denotes the Kullback-Leibler divergence[63]. Larger values of $D_{\text{JS}}(P_{1}\parallel P_{2})$ indicate greater differences between $P_{1}$ and $P_{2}$.
@@
where $P_{\text{rss}}$ is the motif population density distribution of a chemically random solid solution. Larger values of $D_{\text{sro}}$ indicate a greater absolute amount of SRO in the state defined by $P$.
For annealing simulations, $P$ was evaluated using data from all seven snapshots across each of the 65 independent Monte Carlo runs described in section 2.3 (i.e., $7\times 65\times 4{,}394=1{,}999{,}270$ motifs per temperature). For TMP simulations, structures were first denoised using a score-based denoising algorithm[58] for 8 steps. This was followed by identification of the crystal structure using the polyhedral template matching (PTM) algorithm with a root-mean-square deviation cutoff of 0.1. Only atoms identified as bcc were considered for constructing $P$. In both cases, $P$ was generated by randomly sampling subsets of 1,900,000 motifs to ensure consistency across datasets.
To determine whether SRO states obtained from TMP are distinguishable from equilibrium configurations, we adopted the methodology developed in ref. [1]. In this approach, the Jensen-Shannon divergence (eq. 2) between a nonequilibrium motif distribution $P$ (obtained from TMP simulations) and all equilibrium distributions $P_{\text{eq}}(T)$ (obtained from annealing simulations) are computed. The minimum value among these, denoted $D_{\text{eff}}$, represents the distance of a nonequilibrium state from its closest equilibrium counterpart. States with $D_{\text{eff}}$ that fall within their expected equilibrium baseline are classified as quasi-equilibrium states, whereas those that deviate are considered far-from-equilibrium states.
Additional quantification of SRO was obtained using the widely employed Warren-Cowley (WC) parameters [64], defined for each atomic pair $AB$ as:
@@ 3.1. Thermomechanical processing-induced SRO
The stress-strain relationship for a representative processing condition ($T=300\,\text{K}$ and $\dot{\epsilon}=10^{9}\,\text{s}^{-1}$) of the TiTaVW TMP simulations is shown in fig. 1b. The stress-strain curve begins with a clear elastic region, where stress increases linearly with strain up to the yield stress, at which point the four initially seeded dislocation loops begin to move, initiating plastic deformation. Soon after yielding, the system reaches a plateau in which the flow stress fluctuates around a steady mean (i.e., steady-state flow stress). Aside from brief transients caused by changes in the loading direction, the steady flow stress remains consistent throughout the seven deformation cycles. A similar trend is observed in the evolution of dislocation density (fig. 1c), where the steady-state dislocation density $\rho^{\infty}$ is on the order of $10^{17}\,\text{m}^{-2}$. This value is consistent with previous molecular dynamics simulation results[56, 55] and with theoretical estimates of the limiting dislocation density[65]. The steady behavior in both flow stress and dislocation density is observed across all temperature and strain rate conditions considered in here, suggesting that it is a fundamental aspect of the alloy’s thermomechanical response, when dislocations serve as the primary carriers of plasticity[56].
@@
To examine how the SRO of the material evolves under such deformation, we monitored $D_{\text{sro}}$ (eq. 3) as a function of strain, as illustrated in fig. 1d. With plastic deformation, the initially chemically random system forms SRO, evidenced by a rapid increase in $D_{\text{sro}}$ with strain. Remarkably, even at a strain of $0.3$, substantial SRO has already formed, indicating that modest levels of plastic strain — typical in many materials processing conditions — are sufficient to induce SRO in random solid solutions. This has important implications for interpreting the role of SRO on mechanical response during experiments: since SRO begins to evolve from the very onset of plastic deformation, even modest plastic strain can significantly alter the local chemical environment in the alloy[66]. As a result, any subsequent mechanical testing will reflect a modified SRO state, complicating efforts to isolate direct influence of SRO on mechanical properties.
@@
With continued deformation, the system eventually reaches a steady value of $D_{\text{sro}}$, denoted $D^{\infty}_{\text{sro}}$, that is independent of the initial SRO state before deformation, as demonstrated in ref. [1]. This implies that even if we were to initialize the simulation with a more ordered configuration, such as a state annealed at moderate temperatures, the system would still evolve toward the same $D^{\infty}_{\text{sro}}$ shown in fig. 1d. The evolution of $D_{\text{sro}}$ with strain during TMP, across all processing conditions considered here, is well described by the following equation:
@@
where $D^{\text{o}}_{\text{sro}}$ is the initial amount of SRO, and $\Gamma$ and $\lambda$ are rate parameters governing the SRO evolution kinetics. Note that $D^{\infty}_{\text{sro}}=D_{\text{sro}}(\epsilon\rightarrow\infty)=\Gamma/\lambda$, from to eq. 5. For the specific case shown in fig. 1d ($T=300\,\text{K}$ and $\dot{\epsilon}=10^{9}\,\text{s}^{-1}$), the values of $\Gamma$ and $\lambda$ are $2.02\times 10^{-3}$ bits per unit strain and $1.03$ per unit strain, respectively.
@@
The two terms on the right-hand side of eq. 6 reveal the existence of two competing processes governing the SRO evolution[1]. The first term ($\Gamma$) is positive and represents the rate of SRO creation, arising from the inherent chemical bias associated with dislocation motion. Driven by applied mechanical forces, dislocations preferentially traverse through low-energy pathways. The energetics of these pathways are dictated by the local chemical motifs near the dislocation core[30, 1]. As a result, dislocation motion in metallic alloys is chemically selective and promotes the formation of SRO by favoring certain motifs over others. The second term ($\lambda D_{\text{sro}}$) is negative and proportional to the amount of SRO present in the system; it accounts for the destruction of SRO due to the randomizing effects of the mechanically-biased component of dislocation motion. Together, the SRO creation rate $\Gamma$ and annihilation rate $\lambda$ govern the overall evolution of SRO during TMP.
Figure 2: Effect of temperature and strain rate on steady-state SRO. a) Steady-state SRO ($D^{\infty}_{\text{sro}}$) as a function of temperature and strain rate. b) The steady-state SRO states are of far-from-equilibrium kind[1] and they move further from equilibrium as temperature increases, indicated by increasing $D_{\text{eff}}$ (sec. 2.4) c) Warren-Cowley parameters for different pairs indicate that there is a change in pair preference at moderate temperatures for certain elemental pairs (indicated by the purple box), marking a change in the chemical motifs forming the SRO.
@@ 3.2. Effects of processing conditions on steady-state SRO
Since TMP is a nonequilibrium process, the resulting SRO states are likewise nonequilibrium in nature. We quantify their deviation from equilibrium configurations using $D_{\text{eff}}$ (sec. 2.4), a metric that measures the statistical distance between a steady-state SRO configuration and its corresponding equilibrium counterpart using eq. 2. Larger $D_{\text{eff}}$ values indicate greater deviations from equilibrium. As shown in fig. 2b, $D_{\text{eff}}$ increases monotonically with temperature, implying that thermal activation of dislocation motion drives the system progressively further from equilibrium SRO. In equilibrium states, $D_{\text{eff}}$ values are expected to remain below $1.52\times 10^{-3}$ bits. However, as indicated in fig. 2b, the steady-states of SRO obtained through TMP consistently exceed this threshold, confirming their far-from-equilibrium character. These far-from-equilibrium SRO states, observed during TMP, cannot be accessed via equilibrium pathways such as thermal annealing. As a result, TMP opens a new regime of SRO states beyond the traditional equilibrium space.
@@ 3.3. Effects of processing conditions on SRO evolution rate
Figure 3: Effect of temperature and strain rate on SRO evolution rate. Temperature and strain rate dependence of a) SRO creation rate $\Gamma$, and b) SRO annihilation rate $\lambda$. The shaded purple region highlights the temperature range where both $\Gamma$ and $\lambda$ become sensitive to changes in temperature and strain rate. Error bars (one standard deviation) are smaller than the marker size and therefore not visible.
@@
Figure 4: Dislocation-mediated mechanisms of SRO evolution. a) Dislocation configuration at $300\,\text{K}$ and b) $2300\,\text{K}$ after reaching steady state during thermomechanical processing at a strain rate of $10^{8}\,\text{s}^{-1}$. Grey features indicate dislocation-induced defects originating from cross-kinks (approximately 595 at 300 K and 361 at 2,300 K.) c) Steady-state dislocation density $\rho^{\infty}$ decreases with temperature and strain rate. d) Probability distribution of dislocation character at $\dot{\epsilon}=10^{8}\,\text{s}^{-1}$, where $\theta$ is the angle between the Burgers vector and line direction $\theta$. e) Ratio of edge-to-screw character $R$ (eq. 1) as a function of temperature and strain rate.
@@ 3.4. Effects of processing conditions on SRO evolution mechanism
Having established that processing parameters can modulate the SRO creation and annihilation rates — and consequently, SRO evolution — we now turn to the underlying mechanisms responsible for this control. As explained in section 3.1, SRO formation during TMP occurs due to the chemical bias associated with dislocation motion[1]. Any phenomenon during deformation that enhances this chemical bias promotes SRO creation, while phenomena that counteract it contribute to SRO annihilation.
We first examine the role of steady-state dislocation density $\rho^{\infty}$ in modulating the chemical bias. A dense dislocation network restricts dislocation mobility through elastic interactions, geometrical constraints, and dislocation junctions. These constraints are evident in the dislocation configuration shown in fig. 4a, which corresponds to the processing condition of $T=300\,\text{K}$ and $\dot{\epsilon}=10^{8}\,\text{s}^{-1}$. The dislocations appear highly entangled, forming dense forests that inhibit their motion. In contrast, the dislocation configuration in fig. 4b, corresponding to $T=2300\,\text{K}$ at the same strain rate, contains fewer dislocations, allowing for the formation of longer, more mobile dislocation lines. The grey features visible in both systems represent dislocation-induced defects such as vacancies, interstitials, and defect clusters. These defects originate primarily from the self-pinning of screw dislocations, a consequence of kink-pair nucleation on intersecting glide planes[67, 11]. Such defects are notably more prevalent in the system shown in fig. 4a than in fig. 4b (i.e., 64% more defects at 300 K than 2,300 K), further impeding dislocation motion, especially of screw types. Together, these constraints limit the number of energetically distinct pathways accessible to dislocations, thereby suppressing the manifestation of chemical bias in their motion. As a result, elevated dislocation densities and defect concentrations reduce the extent of SRO formation.
@@
While the restrictive influence of dislocation forests and dislocation-induced defects accounts for much of the observed temperature and strain rate dependence, it does not fully explain the discernible transition in SRO evolution behavior at moderate temperatures (fig. 3) or the weak strain rate sensitivity observed at low temperatures. To resolve this, we examine the character of the dislocations present in the system and its temperature dependence. It is well established that at low temperatures, screw dislocations are considerably less mobile than their edge counterparts[29] — a difference that diminishes as temperature increases. The higher concentration of self-pinning features at lower temperatures (fig. 4a) is itself a manifestation of this behavior. In addition, fig. 4d shows that increasing temperature leads to a significant shift in the overall dislocation character toward screw. This phenomenon can be directly visualized in figs. 4a and 4b, where it is clear that the system deformed at $300\,\text{K}$ possesses a larger fraction of edge dislocations compared to $2300\,\text{K}$. This overall shift toward a screw-dominated dislocation population at elevated temperatures, combined with their increased mobility, is mechanistically significant.
Among dislocation types, screw dislocations are more chemically biased in their motion. This is due to their ability to move via kink-pair formation and cross-slip. These mechanisms allow screw dislocations to sample numerous transition pathways, making their mobility more sensitive to the local chemical environment. In contrast, edge dislocations are confined mostly to fixed glide planes, rendering their motion less dependent on chemistry (i.e., more mechanically driven). This distinction has been observed in ref. [29], where screw dislocation mobility was shown to depend sensitively on the local chemical ordering, whereas edge dislocation mobility remained largely unaffected.

### v2

@@ 2.4. SRO quantification
To quantify chemical SRO in the states obtained from both TMP and annealing simulations, we employed the chemical motif framework introduced in refs. [sheriff_quantifying_2024] and [sheriff2024chemicalmotif]. In this approach, the local chemical environment of an atom was described by a motif $\mathcal{M}_{i}$, consisting of the atom and its first-nearest neighbors. We utilized the Euclidean graph neural network architecture[smidt_euclidean_2021] from ref. [sheriff2024chemicalmotif] to identify all motifs $\mathcal{M}_{i}$ in a system and compute their population density distributions $P(\mathcal{M}_{i})$, hereafter denoted simply as $P$ for brevity. The difference in SRO between two states, characterized by motif population density distributions $P_{1}$ and $P_{2}$, was quantified using the Jensen-Shannon divergence:
@@
where $M=\frac{1}{2}(P_{1}+P_{2})$ and $D_{\text{KL}}$ denotes the Kullback-Leibler divergence[mackay2003information]. Larger values of $D_{\text{JS}}(P_{1}\parallel P_{2})$ indicate greater differences between $P_{1}$ and $P_{2}$.
@@
where $P_{\text{rss}}$ is the exact motif population density distribution of a chemically random solid solution, as derived in ref. [sheriff2024chemicalmotif]. Larger values of $D_{\text{sro}}$ indicate a greater absolute amount of SRO in the state defined by $P$.
For annealing simulations, $P$ was evaluated using data from all seven snapshots across each of the 65 independent Monte Carlo runs described in section 2.3 (i.e., $7\times 65\times 4{,}394=1{,}999{,}270$ motifs per temperature). For TMP simulations, structures were first denoised using a score-based denoising algorithm[hsu_score-based_2024] for 8 steps. This was followed by identification of the crystal structure using the polyhedral template matching (PTM) algorithm with a root-mean-square deviation cutoff of 0.1. Only atoms identified as bcc were considered for constructing $P$. Minimally distorted motifs were mapped to ideal bcc motif geometry to ensure consistent motif classification. In both cases, $P$ was generated by randomly sampling subsets of 1,900,000 motifs to ensure consistency across datasets.
To determine whether SRO states obtained from TMP are distinguishable from equilibrium configurations, we adopted the methodology developed in ref. [islam_nonequilibrium_2025]. In this approach, the Jensen-Shannon divergence (eq. 2) between a nonequilibrium motif distribution $P$ (obtained from TMP simulations) and all equilibrium distributions $P_{\text{eq}}(T)$ (obtained from annealing simulations) are computed. The minimum value among these, denoted $D_{\text{eff}}$, represents the distance of a nonequilibrium state from its closest equilibrium counterpart. States with $D_{\text{eff}}$ that fall within their expected equilibrium baseline are classified as quasi-equilibrium states, whereas those that deviate are considered far-from-equilibrium states (see Supplementary Information section 4 for details).
Additional quantification of SRO was obtained using the widely employed Warren-Cowley (WC) parameters [cowley_approximate_1950], defined for each atomic pair $AB$ as:
@@ 3.1. Thermomechanical processing-induced SRO
The stress-strain relationship for a representative processing condition ($T=300\,\text{K}$ and $\dot{\epsilon}=10^{9}\,\text{s}^{-1}$) of the TiTaVW TMP simulations is shown in fig. 1b. The stress-strain curve begins with a clear elastic region, where stress increases linearly with strain up to the yield stress, at which point the four initially seeded dislocation loops begin to move, initiating plastic deformation. Soon after yielding, the system reaches a plateau in which the flow stress fluctuates around a steady mean (i.e., steady-state flow stress). Aside from brief transients caused by changes in the loading direction, the steady flow stress remains consistent throughout the seven deformation cycles. A similar trend is observed in the evolution of dislocation density (fig. 1c), where the steady-state dislocation density $\rho^{\infty}$ is on the order of $10^{17}\,\text{m}^{-2}$. This value is consistent with previous molecular dynamics simulation results[zepeda-ruiz_probing_2017, zepeda-ruiz_atomistic_2021] and with theoretical estimates of the limiting dislocation density[cotterill_does_1977]. The steady behavior in both flow stress and dislocation density is observed across all temperature and strain rate conditions considered here (see Supplementary Information figs. 2 and 3), suggesting that it is a fundamental aspect of the alloy’s thermomechanical response, when dislocations serve as the primary carriers of plasticity[zepeda-ruiz_probing_2017].
@@
To examine how the SRO of the material evolves under such deformation, we monitored $D_{\text{sro}}$ (eq. 3) as a function of strain, as illustrated in fig. 1d. With plastic deformation, the initially chemically random system forms SRO, evidenced by a rapid increase in $D_{\text{sro}}$ with strain. Remarkably, even at a strain of $0.3$, substantial SRO has already formed, indicating that modest levels of plastic strain — typical in many materials processing conditions — are sufficient to induce SRO in random solid solutions. This has important implications for interpreting the role of SRO on mechanical response during experiments: since SRO begins to evolve from the very onset of plastic deformation, even modest plastic strain can significantly alter the local chemical environment in the alloy[han_ubiquitous_2024]. As a result, any subsequent mechanical testing will reflect a modified SRO state, complicating efforts to isolate direct influence of SRO on mechanical properties.
@@
With continued deformation, the system eventually reaches a steady value of $D_{\text{sro}}$, denoted $D^{\infty}_{\text{sro}}$, that is independent of the initial SRO state before deformation, as demonstrated in ref. [islam_nonequilibrium_2025]. This implies that even if we were to initialize the simulation with a more ordered configuration, such as a state annealed at moderate temperatures, the system would still evolve toward the same $D^{\infty}_{\text{sro}}$ shown in fig. 1d. The evolution of $D_{\text{sro}}$ with strain during TMP, across all processing conditions considered here (see Supplementary Information fig. 4), is well described by the following equation:
@@
where $D^{\text{o}}_{\text{sro}}$ is the initial amount of SRO, and $\Gamma$ and $\lambda$ are rate parameters governing the SRO evolution kinetics. Note that $D^{\infty}_{\text{sro}}=D_{\text{sro}}(\epsilon\rightarrow\infty)=\Gamma/\lambda$, from eq. 5. For the specific case shown in fig. 1d ($T=300\,\text{K}$ and $\dot{\epsilon}=10^{9}\,\text{s}^{-1}$), the values of $\Gamma$ and $\lambda$ are $2.02\times 10^{-3}$ bits per unit strain and $1.03$ per unit strain, respectively.
@@
The two terms on the right-hand side of eq. 6 reveal the existence of two competing processes governing the SRO evolution[islam_nonequilibrium_2025]. The first term ($\Gamma$) is positive and represents the rate of SRO creation, arising from the inherent chemical bias associated with dislocation motion. Driven by applied mechanical forces, dislocations preferentially traverse through low-energy pathways. The energetics of these pathways are dictated by the local chemical motifs near the dislocation core[utt_origin_2022, islam_nonequilibrium_2025]. As a result, dislocation motion in metallic alloys is chemically selective and promotes the formation of SRO by favoring certain motifs over others. The second term ($\lambda D_{\text{sro}}$) is negative and proportional to the amount of SRO present in the system; it accounts for the destruction of SRO due to the randomizing effects of the mechanically-biased component of dislocation motion. Together, the SRO creation rate $\Gamma$ and annihilation rate $\lambda$ govern the overall evolution of SRO during TMP.
Figure 2: Effect of temperature and strain rate on steady-state SRO. a) Steady-state SRO ($D^{\infty}_{\text{sro}}$) as a function of temperature and strain rate. $D^{\infty}_{\text{sro}}$ values were obtained by averaging $D_{\text{sro}}$ from the seventh compression cycle. b) The steady-state SRO states are of far-from-equilibrium kind[islam_nonequilibrium_2025] and they move further from equilibrium as temperature increases, indicated by increasing $D_{\text{eff}}$ (sec. 2.4) c) Warren-Cowley parameters for different pairs indicate that there is a change in pair preference at moderate temperatures for certain elemental pairs (indicated by the purple box), marking a change in the chemical motifs forming the SRO.
@@ 3.2. Effects of processing conditions on steady-state SRO
Since TMP is a nonequilibrium process, the resulting SRO states are likewise nonequilibrium in nature. We quantify their deviation from equilibrium configurations using $D_{\text{eff}}$ (sec. 2.4), a metric that measures the statistical distance between a steady-state SRO configuration and its corresponding equilibrium counterpart using eq. 2. Larger $D_{\text{eff}}$ values indicate greater deviations from equilibrium. As shown in fig. 2b, $D_{\text{eff}}$ increases monotonically with temperature, implying that thermal activation of dislocation motion drives the system progressively further from equilibrium SRO. In equilibrium states, $D_{\text{eff}}$ values are expected to remain below $1.52\times 10^{-3}$ bits (see Supplementary Information section 4). However, as indicated in fig. 2b, the steady-states of SRO obtained through TMP consistently exceed this threshold, confirming their far-from-equilibrium character. These far-from-equilibrium SRO states, observed during TMP, cannot be accessed via equilibrium pathways such as thermal annealing. As a result, TMP opens a new regime of SRO states beyond the traditional equilibrium space.
@@ 3.3. Effects of processing conditions on SRO evolution rate
[…]
smaller than the marker size and therefore not visible.
@@
Figure 4: Dislocation-mediated mechanisms of SRO evolution. a) Dislocation configuration at $300\,\text{K}$ and b) $2300\,\text{K}$ after reaching steady state during thermomechanical processing at a strain rate of $10^{8}\,\text{s}^{-1}$. Grey features indicate dislocation-induced defects originating from cross-kinks (approximately 595 at 300 K and 361 at 2,300 K.) c) Steady-state dislocation density $\rho^{\infty}$ decreases with temperature and strain rate. $\rho^{\infty}$ values were obtained by averaging $\rho$ from the seventh compression cycle. d) Probability distribution of dislocation character at $\dot{\epsilon}=10^{8}\,\text{s}^{-1}$, where $\theta$ is the angle between the Burgers vector and line direction $\theta$. e) Ratio of edge-to-screw character $R$ (eq. 1) as a function of temperature and strain rate.
@@ 3.4. Effects of processing conditions on SRO evolution mechanism
Having established that processing parameters can modulate the SRO creation and annihilation rates — and consequently, SRO evolution — we now turn to the underlying mechanisms responsible for this control. As explained in section 3.1, SRO formation during TMP occurs due to the chemical bias associated with dislocation motion[islam_nonequilibrium_2025]. Any phenomenon during deformation that enhances this chemical bias promotes SRO creation, while phenomena that counteract it contribute to SRO annihilation.
We first examine the role of steady-state dislocation density $\rho^{\infty}$ in modulating the chemical bias. A dense dislocation network restricts dislocation mobility through elastic interactions, geometrical constraints, and dislocation junctions. These constraints are evident in the dislocation configuration shown in fig. 4a, which corresponds to the processing condition of $T=300\,\text{K}$ and $\dot{\epsilon}=10^{8}\,\text{s}^{-1}$. The dislocations appear highly entangled, forming dense forests that inhibit their motion. In contrast, the dislocation configuration in fig. 4b, corresponding to $T=2300\,\text{K}$ at the same strain rate, contains fewer dislocations, allowing for the formation of longer, more mobile dislocation lines. The grey features visible in both systems represent dislocation-induced defects such as vacancies, interstitials, and defect clusters. These defects originate primarily from the self-pinning of screw dislocations, a consequence of kink-pair nucleation on intersecting glide planes[zhou_cross-kinks_2021, marian_dynamic_2004]. Such defects are notably more prevalent in the system shown in fig. 4a than in fig. 4b (i.e., 64% more defects at 300 K than 2,300 K), further impeding dislocation motion, especially of screw types. Together, these constraints limit the number of energetically distinct pathways accessible to dislocations, thereby suppressing the manifestation of chemical bias in their motion. As a result, elevated dislocation densities and defect concentrations reduce the extent of SRO formation.
@@
While the restrictive influence of dislocation forests and dislocation-induced defects accounts for much of the observed temperature and strain rate dependence, it does not fully explain the discernible transition in SRO evolution behavior at moderate temperatures (fig. 3) or the weak strain rate sensitivity observed at low temperatures. To resolve this, we examine the character of the dislocations present in the system and its temperature dependence. It is well established that at low temperatures, screw dislocations are considerably less mobile than their edge counterparts[yin_atomistic_2021] — a difference that diminishes as temperature increases. The higher concentration of self-pinning features at lower temperatures (fig. 4a) is itself a manifestation of this behavior. In addition, fig. 4d shows that increasing temperature leads to a significant shift in the overall dislocation character toward screw. This phenomenon can be directly visualized in figs. 4a and 4b, where it is clear that the system deformed at $300\,\text{K}$ possesses a larger fraction of edge dislocations compared to $2300\,\text{K}$. This overall shift toward a screw-dominated dislocation population at elevated temperatures, combined with their increased mobility, is mechanistically significant.
Among dislocation types, screw dislocations are more chemically biased in their motion. This is due to their ability to move via kink-pair formation and cross-slip. These mechanisms allow screw dislocations to sample numerous transition pathways, making their mobility more sensitive to the local chemical environment. In contrast, edge dislocations are confined mostly to fixed glide planes, rendering their motion less dependent on chemistry (i.e., more mechanically driven). This distinction has been observed in ref. [yin_atomistic_2021], where screw dislocation mobility was shown to depend sensitively on the local chemical ordering, whereas edge dislocation mobility remained largely unaffected.